OpenStudy (anonymous):

How do I prove this an Identiy?: ((cos(a+b))/(cosa +sinb))=((cosa - sinb)/(cos(b-a))) Please help. any help would be appreciated

OpenStudy (anonymous):

what is cos(a+b) ? and almost the same, what is cos(b-a) ?

OpenStudy (anonymous):

cos(a+b)= (cosa)(cosb)-(sina)(sinb) cos(b-a)= (cosb)(cosa)+(sinb)(sina)

OpenStudy (anonymous):

my proof isnt going where I thought it would, sorry

OpenStudy (anonymous):

thanks for trying

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

(cos(a)+sin(b))x(cos(a)-sin(b))=cos^2(a)-sin^2(b) and then use the cosxcosy=1/2 (cos(x-y)+cos(x+y)) identity to figure out what the other end comes out. I can't guarantee it to work, but I think you should give it a try.

OpenStudy (anonymous):

I couldn't do it, but if you can simplify ((cos(a+b))/(cosa +sinb)) - ((cosa - sinb)/(cos(b-a))) down to 0, you will have your proof, just not in the right form. It is possible, but not for me (yet). Again, sorry I couldn't be more help.